OpenStudy (anonymous):

A circular pie with a diameter of 11.2 inches is cut into 8 equal slices. What is the approximate area of one slice?

OpenStudy (rulnick):

area of slice = (1/8) * area of pie = (1/8) * pi (11.2/2)^2

OpenStudy (rulnick):

... about 12.32 in^2.